OpenStudy (anonymous):

Write a word problem describing the time it takes to complete this activity individually and together. For example, if John takes 2 hours to mow his lawn and it takes his sister, Maria, 4 hours to mow the same lawn, how long would it take John and Maria to mow the lawn together? DO NOT USE THE EXAMPLE i need to make my own and i decided to do building a treehouse, but i don't know how to write it out

OpenStudy (anonymous):

@jim_thompson5910 @myininaya

OpenStudy (anonymous):

"Leah and Caleb both built their own tree houses that were of equal size. It took Caleb 5 hours to complete his tree house, and it took Leah 8 hours. Since both tree houses are of equal size, how much time would elapse if Caleb and Leah worked together to build a tree house?"

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

can you help me with a couple more that build off this?

OpenStudy (anonymous):

I'll gladly try. I'm not a math genius myself, but I can see if I can help.

OpenStudy (anonymous):

thanks. so you have to write a rational based on the word problem

OpenStudy (anonymous):

for example, for the lawn mowing it would be \[\frac{ 1 }{ 2 }+\frac{ 1 }{ 4 }= \frac{ 1 }{ x}\]

OpenStudy (anonymous):

Alright, let me take a look.

OpenStudy (anonymous):

I'm not quite sure about this one. I'm sorry, I just don't want to give you wrong answers.

OpenStudy (anonymous):

its ok man ! no problem